OpenStudy (noelgreco):

You are given a circular piece of cardboard, and asked to cut out a sector so that when the cut edges are brought together a cone of maximum volume is formed. What central angle maximizes the volume of the cone? a.) Decimal approximation b.) Exact answer

OpenStudy (datanewb):

I get that the volume V relates to the central angle in this way:\[V=\frac{\pi}{3}sin^2(\alpha)cos(\alpha)\] ... but taking the derivative has not allowed me to deduce the maximum. Is that equation correct? If so, how does one find the root of \[\frac{\pi}{3}(2sin(\alpha) cos^2(\alpha) - sin^3(\alpha)) \]

OpenStudy (noelgreco):

I didn't use trigonometry in my calculation, but in graphing your formula for V I got a maximum at a different value.

OpenStudy (noelgreco):

\[\Theta=\frac{ 6\pi -2\pi \sqrt{6} }{ 3 }\], or 1.15299 radians \[66.06^{o}\]

OpenStudy (datanewb):

 oops I misunderstood what a central angle was! Let me recompute!